OpenStudy (anonymous):

(f(10+h)-f(10))/h can be written as: -1/A(B+h) Where are A and B constants? F(X)=1/(x+2)

OpenStudy (anonymous):

If this helps: A=? B=?

OpenStudy (anonymous):

F(X)=1/(x+2) F(x+h)=1/(x+2+h) (F(x+h)-F(x))/h=1/h [1/(x+2+h)-1/(x+2)]=(x+2-x-2-h)/ h(x+2)(x+2+h)=-1/(x+2)(x+2+h)